# theoTheorem

### v2

@@ Abstract
Laser phase noise is one of the main sources of decoherence in driven Rydberg systems with neutral atoms in tweezer arrays. While the effect of phase noise in the regimes of isolated qubits and few-qubit gate protocols has been studied extensively, there are open questions about the effects of this noise on many-body systems. In many scenarios, the effects of noise cannot simply be described by an increase in the energy or temperature of the system, leading to non-trivial changes in the state and relevant correlations. In this work, we use stochastic sampling to simulate laser phase noise based on experimentally relevant spectral densities. We explore the impact of this noise on adiabatic state preparation in a one-dimensional system, discussing the interplay between heating and interactions during dynamics. We find that for certain adiabatic processes, the noise can be seen to approximately thermalize in terms of its effects on relevant correlation functions.
@@ I Introduction
Progress on the scalability of optical trapping and coherent control have made neutral atom arrays an exciting platform for quantum computation and simulation [1, 2, 3, 4]. In these systems, strong inter-particle interactions between individually trapped atoms are realized by exciting atoms to high-energy Rydberg levels, giving rise to a blockade effect [5] that can be leveraged for the creation of quantum gates [6, 7, 8], as well as to a range of spin Hamiltonians [9, 34]. Applications of these systems include the exploration of annealing processes relevant for optimization [10, 11], realizing topological models [12, 13], spin liquid phases [14, 15, 16], quantum many-body scars [17] and information scrambling [18, 19, 20]. Much of the progress in this area in the past years came from reductions in laser phase noise, which crucially affects excitations to Rydberg levels and directly generates dephasing [21, 22, 23, 24]. The effects of this laser phase noise have mainly been quantified in the context of single qubit operations. In simulation of quantum dynamics, an important question is how this single-particle heating affects the overall behavior of a many-body system. This connects to fundamental questions on thermalization in closed quantum systems [25, 26, 27, 28, 29], specifically whether the heating due to laser phase noise results in correlation functions that would be predicted by the increase in energy and temperature.
In this work, we quantify the effects of laser phase noise on a many-body spin systems as realized in tweezer arrays. The adiabatic preparation of many-body ground states provides an important example in this context, with the procedure used to prepare interesting many-body states [30, 31, 32, 9], as well as in quantum annealing [33, 34, 35, 36]. We simulate time evolution of an adiabatic state preparation of our antiferromagnetic spin chain, applying laser phase noise that has been sampled from experimentally realistic diode laser data, similar to phase noise measured, e.g., in Ref. [21]. We extract features that are relevant to general noise spectra, quantifying the reduction of fidelity and changes in correlation functions that result from adding phase noise with varying strengths and frequency profiles relative to central frequencies associated with the system dynamics. We analyze in detail the couplings to excited states generated by noise, and also show to what extent (despite the complicated many-body spectra) sensitivity to noise can be reduced through the choice of the relative frequencies. We then also ask to what extent we might have expected the results we obtain for correlations based on thermalization of the states. We evaluate the long-time average expectation values of relevant observables, comparing them to predictions by a thermal equilibrium state of the same energy to see how closely the relaxation of energy in the system approaches thermalization.
@@
Figure 1: Parameter space plot for the Rydberg Hamiltonian seen in Eq. (II) for a chain of N=11 atoms. (a) A schematic diagram showing crystalline ordering of ground (blue) and Rydberg excited (red) atoms for initial and target many-body state of adiabatic state preparation. (b) Total fraction of excited Rydberg atoms $N_{R}/N$ in the ground state plotted against laser parameters $\Omega/\delta$, from no excitations (blue), to half filled (white), to fully excited (red).The path for the diabatic state preparation studied is provided with color gradient from white to green tracking ground state energy gap. All parameters are provided in units interaction strength $V_{dd}$.
@@ II The physical system
The simulations discussed in this paper are carried out in a one-dimensional Rydberg spin Hamiltonian that takes the form
@@
where the first term drives excitations in the system at a Rabi frequency $\Omega(t)$ modulated by a phase noise signal $\phi(t)$ that introduces a small error at every time step. The parameter $\delta(t)$ models the global laser detuning, and the operator $n_{k}\equiv\sigma_{z}^{k}-\hat{I}$ counts the number of Rydberg excitations in site $k$. The terms $\delta(t)$ and $\Omega(t)$ are tuneable in an experimental setting and are governed by the laser frequency, amplitude, and phase stability. Finally, we define the interaction Hamiltonian $\hat{H}_{\text{int}}$ explicitly to use it as an observable
@@
introducing a dipole-dipole potential $V_{dd}$ mediated by $\hat{n}_{k}\hat{n}_{l}$ interactions that enforce a Rydberg blockade penalty on neighboring excitations. We work in the dipole dipole regime in which long range interactions decay as a function of distance between sites $d^{3}=|k-l|^{3}$, where $k$ and $l$ are site indices and the lattice constant is absorbed into $V_{dd}$. Open boundary conditions are considered for the one-dimensional chain. For computational simplicity, we define a relative number to which we set the interaction strength $V_{dd}/2\pi=1/2\pi$  MHz, and use this to scale all units in the system with $\hbar=1$ also assumed. Hence, both $\Omega$ and $\delta$ are expressed in units of this relative $V_{dd}$, while time is given in units of $V_{dd}^{-1}$. The choice allows us to study many body dynamics in a similar frequency region as the sampled noise, and the results in this paper can be generalized to different frequency regimes. Note that for $\Omega=0$ the Hamiltonian in Eq. (II) is reminiscent of the Ising model for magnetic dipoles. In such a regime, the competition between $\delta$ and $V_{dd}$ leads to an excitation ladder in which ground states with increasing number of evenly spaced excitations are favored [30, 32, 37]. These crystalline ground states exhibit large energy gaps that make them robust to small increases in $\Omega$, leading to regions in the $\Omega/\delta$ parameter space where only one crystalline state is represented in the ground state. In Fig. 1 we show Rydberg excitations along the spin chain as a function of $\Omega/\delta$ parameters, clearly displaying the formation of a lobe with a stable ground state with crystalline $Z_{2}$ order of excitations. How close the final prepared many-body ground state $|\psi_{gr}\rangle$ is to a perfect $Z_{2}$ ordering can always be tested by computing the expectation value of the order parameter observable
@@
where we check anti-ferromagnetism over all different lattice sites $k$ and $l$. The goal of the procedure is to take a ground state with no Rydberg excitations to a ground state with $Z_{2}$ ordered excitations and expectation value $\langle\hat{O}_{Z_{2}}\rangle=1$, while maximizing the ground state energy gap to mitigate diabatic excitation. Due to the open boundary conditions, atoms found on the edges of the chain experience less of the effect of Rydberg blockade and are therefore more likely to be Rydberg excited than atoms within the chain. This means that energetically favorable $Z_{2}$ ordering of excitations can be identified uniquely in chains with an odd number of lattice sites, but an even number of lattice sites results in a central domain wall after which ordering flips [9]. To simplify notation and discussions, here we will focus on odd numbers of spins. Also shown in Fig. 1 is the path of the adiabatic state preparation simulated in this paper, broken up into a series of three linear ramps that tune $\Omega$ and $\delta$ independently. The next section will introduce this state preparation in more detail as well as providing results after introducing phase noise to the system.
@@ III Phase noise in adiabatic state preparation
$|\langle E_{m}(t)|\dot{H}(t)|E_{n}(t)\rangle|\ll|E_{m}(t)-E_{n}(t)|,$ (4)
where $|E_{m}(t)\rangle$ and $|E_{n}(t)\rangle$ are eigenstates with energies $E_{m}(t)$ and $E_{n}(t)$ in the instantaneous Hamiltonian with an energy gap that we do not wish to cross. Typically, the procedure is done starting from a state which is easy to prepare experimentally to a target state which is harder to access. The challenge in adiabatic state preparation is thus to maximize the energy gap to the first excited state $\Delta E_{gr}$ and to decrease the variation rate if $\Delta E_{gr}$ becomes small enough to result in unwanted excitation of higher energy eigenstates.
@@ III.1 The adiabatic ramp
Figure 3: Realistic noise power spectrum used to generate independent phase noise time realizations implemented in adiabatic time evolution. (a) shows the phase noise spectrum $S_{\phi}$ (green) which has been converted from a frequency spectrum $S_{\nu}$ (red) used in noise generation, with dashed line marking system interaction strength $V_{dd}$. The relevant conversion equation is $S_{\phi}=\nu^{2}S_{\nu}$. Also provided in (b) are three independent noise realizations generated using $S_{\phi}$ (see Appendix A). The averaged noise behavior after 100 independently generated noise realizations is shown in blue.
Both laser parameters $\delta$ and $\Omega$ in the Hamiltonian in Eq. (II) are time dependent and can be tuned independently. Starting from a Hamiltonian with large negative detuning in a ground state with no Rydberg excitations, Fig. 2(a) shows the preparation of the $Z_{2}$ target state using a three-step adiabatic ramp originally proposed in [31]. The initial Hamiltonian is far detuned at laser parameters $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=-3$ where a large ground state gap allows for easy experimental preparation. The target Hamiltonian occurs at $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=1.1$ with a much smaller gap and a ground state with $Z_{2}$ excitation ordering. Figures 2(b,c) show the change in the ground state fidelity $|\langle\psi|\psi_{gr}\rangle|^{2}$ and ground state energy gap $\Delta E_{gr}$ of such an adiabatic state preparation for a chain of 11 atoms. Since the evolution time is finite and thus not perfectly adiabatic, there is a small amount of excitation throughout the process leading to a drop in overall fidelity. As the target Hamiltonian is approached $\Delta E_{gr}$ decreases considerably, making excitation of higher energy many-body states more probable, thus reducing the final fidelity. Most unwanted excitations occur during the critical region in the final step of preparation meaning that the variation rate of the Hamiltonian must be slowed to maintain fidelity with the ground state. Given the resource intensive nature of many-body time evolution we restrict ourselves to simulating the final section of the ramp as it is the most critical, testing it for different variation rates.
@@ III.2 Introducing phase noise
Figure 4: Final fidelity (a) and total energy per particle relative to the ground state (b) for different ramp times $T_{3}$ of the final step in adiabatic state preparation described in Fig 2 for various system sizes $N$. Black line shows results for 11 sites with no phase noise, while all other results have phase noise. Optimal values of $T_{3}$ for a given $N$ which result in highest fidelity are highlighted in red. Error bars show standard error after averaging over 100 preparations with independent laser phase noise realisations. These results have been calculated using tensor network simulations using the iTensors package in Julia [38].
Figure 5: Final fidelity of prepared ground state with fixed duration $T_{3}V_{dd}=90$ and different system sizes $N$, applying a variety of scaled phase noise. (a) Final fidelity after direct scaling of applied phase noise $\phi(t)\rightarrow m\phi(t)$. (b) Final fidelity after re-scaling frequency grid of sampled spectral density by a factor of $\kappa$, moving the the noise profile peaked at $3V_{dd}\approx 2\pi\times 480$  kHz (see Fig. 3) to higher or lower frequencies. Modified spectra are renormalized to ensure that total noise power remains constant. Dashed line shows scaling for which noise is most detrimental to fidelity. Error bars show standard error after averaging 100 preparations with laser phase independent noise realizations.
Figure 6: Matrix elements of the many-body operator $\hat{\sigma}_{z}=\sum_{k}^{N}\hat{\sigma}^{k}_{z}$ for the initial Hamiltonian (a), with laser parameters $\Omega/\text{V}_{dd}=1.0$ and $\delta/\text{V}_{dd}=1.1$, as well as the final Hamiltonian (b), with parameters $\Omega/\text{V}_{dd}=0.1$ and $\delta/\text{V}_{dd}=1.1$, of the third stage of the adiabatic state preparation initialized in the ground state $|E_{0}\rangle$ for 11 sites. Dashed lines mark energies of four energy eigenstates chosen for the analysis due to their high occupancy after a time evolution of the two time independent Hamiltonians with phase noise (see appendix B Fig. 11). Points for the same color represent off-diagonal matrix elements.
@@
When noise is added to the transverse field, the resulting dephasing of the dynamics accumulates and, over time, causes unwanted excitations. Noise is added to the system with $\phi(t)$, which is implemented as a phase term modulating $\Omega(t)$ in Eq. (II). Realizations of $\phi(t)$ are generated using an algorithm that stochastically samples a power spectrum chosen to be typical for Rydberg lasers. Fig. 3 provides the original frequency power spectrum of laser noise for a diode laser, and transforms it into a phase power spectrum that can then be used to generate unique discrete noise signals. Details on the simulation of realistic phase noise and the excitation dynamics that lead to the results in this paper are provided in the Appendices A and B.
Fig. 4 shows the final fidelity of preparing the target ground state and energy per particle for a variety of different durations of the final step of the adiabatic state preparation $T_{3}$ with added phase noise. The initial fidelity is set to 1, assuming perfect ground state preparation for steps 1 and 2. As expected, when no laser noise is present we observe constant but diminishing improvement in fidelity with increasing $T_{3}$, as the slower variation rate approaches the adiabatic limit. The addition of phase noise results in an overall drop in efficacy of the preparation, and when taken to long evolution times fidelity begins to decrease. It is important to note here that if the two initial steps in the adiabatic ramp were included in the simulation the accumulated phase noise across the full procedure would be much greater, leading to further degradation of the ground state. There is an emergent competition between diabatic and phase noise excitation, with both sources of noise exhibiting different excitation profiles and antagonistic relationships to evolution time. This leads to an emergence of an optimal $T_{3}$ at which total energy added to the system is at a minimum, corresponding to the largest achievable fidelity in state preparation. This optimal $T_{3}$ can be seen in Fig. 4 to increase with system size as the variation rate must reduce considerably to mitigate diabatic excitation irrespective of the presence of noise. This is consistent with recent analysis [39], predicting exceedingly long adiabatic ramps for larger system sizes.
Fig. 5 demonstrates the impact that laser phase noise has on the performance of the state preparation as both strength and relative frequency profile of the noise are varied for a fixed ramp duration of $T_{3}V_{dd}=90$. In the case of noise strength, Fig. 5(a) introduces a factor $m$ that linearly scales the phase noise to $\phi(t)\rightarrow m\phi(t)$. The results demonstrate the susceptibility of such adiabatic protocols to phase noise, as we observe a dramatic drop in fidelity with the target ground state in a single order of magnitude amplification. In contrast, Fig. 5(b) shows a modest improvement in fidelity as the sampled power spectrum used for noise generation is modified to have the peak of the noise moved further away from the frequencies of $\Omega$ used during the adiabatic ramp. The modification involves a linear scale factor $\kappa$ applied to the frequency grid of the original spectrum seen in the $x$-axis of Fig. 3(a), thus maintaining the relative shape of the spectrum while controlling the frequency region where the noise power is most prominent. Care was taken to ensure that the total power of the original noise spectral density is preserved, by renormalizing the modified spectrum such that its integral across the relevant frequencies is the same as that of the original. Testing a constant ramp for various modified noise spectra with peaks both above and below the original spectrum, Fig. 5 shows that moving the peak power of the phase noise away from the frequencies taken during the preparation by the optical drive $\Omega(t)$ leads to improved fidelity. Phase noise made up of frequencies that are too high or too low relative the dynamics of the system inhibit the ability of the noise to transfer power into the system, thereby causing less excitation. This suggests that even if phase noise is present in a laser, the effects of phase noise on many-body experiments can be effectively mitigated by moving dynamics far away from affected frequencies as was shown for single atoms and two qubit gates in [21, 22].

### v3

# arXiv:2505.04595v3[quant-ph] 02 Mar 2026

@@ Abstract
Laser phase noise is one of the main sources of decoherence in driven Rydberg systems with neutral atoms in tweezer arrays. While the effect of phase noise in the regimes of isolated qubits and few-qubit gate protocols has been studied extensively, there are open questions about the effects of this noise on many-body systems. In many scenarios, the effects of noise cannot simply be described by an increase in the energy or temperature of the system, leading to non-trivial changes in the state and relevant correlations. In this work, we use stochastic sampling to simulate laser phase noise based on experimentally relevant spectral densities. We explore the impact of this noise on adiabatic state preparation in a one-dimensional system, discussing the interplay between dephasing and diabatic excitation during dynamics. Moreover, we investigate the approximate thermalization of dephasing energy in an interaction regime resonant with the noise, showing convergence to canonical ensemble predictions for two relevant observables.
@@ I Introduction
Progress on the scalability of optical trapping and coherent control have made neutral atom arrays an exciting platform for quantum computation and simulation [52, 3, 24, 43]. In these systems, strong inter-particle interactions between individually trapped atoms are realized by exciting atoms to high-energy Rydberg levels, giving rise to a blockade effect [28] that can be leveraged for the creation of quantum gates [41, 17, Pelegrí_2022], as well as to a range of spin Hamiltonians [4, 16]. Applications of these systems include the exploration of annealing processes relevant for optimization [59, 44], realizing topological models [13, 62], spin liquid phases [57, 21, 36], quantum many-body scars [6] and information scrambling [23, 40, 5]. Much of the recent progress in this area came from reductions in laser phase noise, which crucially affects excitations to Rydberg levels and directly generates dephasing [12, 39, 10, 27]. The effects of this laser phase noise have mainly been characterized in the context of single qubit operations. However, in simulation of quantum dynamics, an important question this paper asks is how this noise affects the overall behavior of a many-body system. This connects to fundamental questions about thermalization in closed quantum systems [51, 8, 54, 9, 15], specifically whether heating due to laser phase noise results in correlation functions predicted by the increase in energy and temperature.
In this work, we quantify the effects of laser phase noise on many-body atomic spin systems as realized in tweezer arrays. The adiabatic preparation of many-body ground states provides a simple but relevant test bed for our study of laser phase noise, used both to prepare interesting many-body states [48, 53, 55, 4], as well as in quantum annealing [29, 16, 22, 33]. We simulate the time evolution of an adiabatic state preparation of our antiferromagnetic spin chain, applying laser phase noise that has been sampled from experimentally realistic diode laser data, similar to phase noise measured, e.g., in Ref. [12]. Furthermore, to avoid complicating the discussion and obscuring the effects of noise, we work with a linear adiabatic ramp in the dipole interaction regime, typical of early theoretical [53] and experimental demonstrations [48] of this adiabatic protocol. We extract features that are relevant to general noise spectra, using the loss of ground state fidelity to quantify the competition between diabatic and dephasing excitation that occurs when we change laser parameter variation rates, as well as showing the effect of changing noise strength, and interaction strength between sites. We supplement these results with a nuanced look into excitation dynamics in our noise, providing an analysis on available transition elements that occur at different stages of the protocol, and how these lead to distinct excitation patterns in time independent Hamiltonian evolutions. Next, we ask to what extent can the energy added to the system by phase noise lead to apparent thermalization in two relevant observables. To answer this question, we work in an interaction regime that is considerably more resonant with the applied noise, allowing us to clearly observe mechanisms through which dephasing excites a many-body system. We evaluate the long-time average expectation values of relevant observables, comparing them to predictions by a thermal equilibrium state of the same energy to see how closely the relaxation of energy resembles thermalization.
@@
[…]
@@ II The physical system
The simulations discussed in this paper are carried out in a one-dimensional Rydberg spin Hamiltonian of the form
@@
[…]
@@
introducing a dipole-dipole interaction strength $V_{dd}$ mediated by $\hat{n}_{k}\hat{n}_{l}$ interactions that enforce a dipolar blockade penalty on neighboring excitations. We work in the dipole dipole interaction regime in which Rydberg states must maintain a permanent dipole moment that is typically achieved through strong electric fields [2, 7, 11], and long range interactions decay as a function of distance between sites $d^{3}=|k-l|^{3}$, where $k$ and $l$ are site indices and the lattice constant is absorbed into $V_{dd}$. Open boundary conditions are considered for the one-dimensional chain, for which atoms found on the edges of the chain experience less of the effect of Rydberg blockade and are therefore more likely to be Rydberg excited than atoms within the chain. This means that energetically favorable $Z_{2}$ ordering of excitations can be identified uniquely in chains with an odd number of lattice sites, but an even number of lattice sites results in a central domain wall after which ordering flips [4]. To simplify notation and discussion, here we will focus on odd numbers of spins. Furthermore, while the noiseless protocol can be generalized across different frequency regimes, for computational simplicity we define an experimentally reasonable relative number to which we set the interaction strength $V_{dd}/2\pi=10$ MHz [49, 4], and use this to scale all units in the system with $\hbar=1$ also assumed. Hence, both $\Omega$ and $\delta$ are expressed in units of this relative $V_{dd}$, while time is given in units of $V_{dd}^{-1}$. Note that for $\Omega=0$ the Hamiltonian in Eq. (II) is reminiscent of the Ising model for magnetic dipoles. In such a regime, the competition between $\delta$ and $V_{dd}$ leads to an excitation ladder in which ground states with increasing number of evenly spaced excitations are favored [48, 55, 47]. These crystalline ground states exhibit large energy gaps that make them robust to small increases in $\Omega$, leading to regions in the $\Omega/\delta$ parameter space where the ground state is dominated by a single crystalline state. Figure 1(a) shows a diagram of the ground state of the initial and final Hamiltonian for the three part adiabatic state preparation discussed in Section III, while in Figure 1(b) we show the path that this protocol takes along the $\Omega$ and $\delta$ parameter plot. The figure also shows Rydberg excitations along the spin chain as a function of $\Omega$ and $\delta$ parameters, clearly displaying the formation of a lobe where a stable ground state with crystalline $Z_{2}$ order of excitations persists. How close the final prepared many-body ground state $|\psi_{gr}\rangle$ is to a perfect $Z_{2}$ ordering can be tested by computing the expectation value of the order parameter
@@
where we check anti-ferromagnetism over all different lattice sites $k$ and $l$. The goal of the procedure is thus to take a ground state with no Rydberg excitations to a ground state with $Z_{2}$ ordered excitations and expectation value $\langle\hat{O}_{Z_{2}}\rangle=1$, while maximizing the ground state energy gap to mitigate diabatic excitation. Finally in Figure 1(c) we also provide a surface plot of the ground state energy gap $\Delta E_{gr}/V_{dd}$ again as a function of $\Omega/V_{dd}$ and $\delta/V_{dd}>0$, along with a highlighted path taken through this region by the adiabatic protocol. The surface plot shows the emergence of a saddle point as $\Delta E_{gr}/V_{dd}$ drops at the transition point to a $Z_{2}$ ordered ground state, resulting in a critical region where low $\Delta E_{gr}/V_{dd}$ could result in increased excitation during a linear adiabatic ramp. The next section will introduce this state preparation in more detail as well as providing results after introducing phase noise to the system.
@@ III Phase noise in adiabatic state preparation
$|\langle m(t)|\dot{H}(t)|n(t)\rangle|\ll|E_{m}(t)-E_{n}(t)|^{2},$ (4)
where we maintain $\hbar=0$ and with $|m(t)\rangle$ and $|n(t)\rangle$ being eigenstates corresponding with energies $E_{m}(t)$ and $E_{n}(t)$ in the instantaneous Hamiltonian with an energy gap that we do not wish to cross. Typically, the procedure is done starting from a state which is easy to prepare experimentally to a target state which is harder to access. The challenge in adiabatic state preparation is thus to maximize the energy gap to the first excited state $\Delta E_{gr}$ and to decrease the variation rate if $\Delta E_{gr}$ becomes small enough to result in unwanted excitation of higher energy eigenstates.
@@ III.1 The adiabatic ramp
Figure 3: Realistic noise power spectrum used to generate independent phase noise time realizations implemented in adiabatic time evolution. (a) The phase noise spectrum $S_{\phi}$ (green) which has been converted from a frequency spectrum $S_{\nu}$ (red) used in noise generation. The relevant conversion equation is $S_{\phi}=\nu^{2}S_{\nu}$. (b) Three independent noise realizations generated using $S_{\phi}$ (see Appendix A) provided in units of $tV_{dd}$ where $V_{dd}/2\pi=10$MHz. The averaged noise behavior after 100 independently generated noise realizations is shown in blue.
Figure 4: Final fidelity of prepared state $|\psi(T)\rangle$ with the instantaneous ground state $|\psi_{gr}(T)\rangle$ at the end of the ramp, and final energy $E_{T}$ relative to the ground state energy $E_{gr}=\langle\psi_{gr}|H_{R}|\psi_{gr}\rangle$ after simulations using different evolution times $T_{2}$ for stage 2 (a,b), and $T_{3}$ for stage 3 (c,d) in the adiabatic state preparation described in Figure 1 with added laser phase noise. Colored lines show results for noisy state preparation with a variety of system sizes, while the black line provides a reference for 11 sites and no added noise. Optimal values of $T_{2}$ resulting in highest fidelity for a given NN are highlighted in larger magenta circles in (a), with their final states used for $T_{3}$ analysis in (c,d). Each data point has error bars showing associated standard error after averaging over 100 preparations with unique laser phase noise signals generated from the same power spectrum in Figure 3. These results have been calculated using tensor network simulations using the iTensor package in Julia [18], with time evolution achieved using the TDVP algorithm with a maximum bond dimension of 100.
Both laser parameters $\delta$ and $\Omega$ in the Hamiltonian in Eq. (II) are time dependent and can be tuned independently. Starting from a Hamiltonian with large negative detuning in a ground state with no Rydberg excitations, Figure 2(a) shows the preparation of the $Z_{2}$ target state using a variation of the three-step adiabatic ramp originally proposed in [53]. The initial Hamiltonian is far detuned at laser parameters $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=-3$ where a large ground state gap allows for reliable experimental preparation. The target Hamiltonian occurs at $\Omega/V_{dd}=0.1$ and $\delta/V_{dd}=1.1$ with a much smaller gap and a ground state with $Z_{2}$ excitation ordering. Figure 2(a) plots the linear tuning of the two laser parameters across all three stages of the protocol, while Figures 2(b,c) show the change in ground state fidelity $|\langle\psi|\psi_{gr}\rangle|^{2}$ and ground state energy gap $\Delta E_{gr}$ of such an adiabatic state preparation for a chain of 11 atoms. Since the evolution time is finite and thus not perfectly adiabatic, there is a small amount of excitation throughout the process leading to a drop in overall fidelity. At the transition between stages 2 and 3 $\Delta E_{gr}$ decreases considerably as shown in Figure 1(c), dramatically increasing the likelihood of exciting higher many-body energy states. Most unwanted excitations occur during this critical region in the adiabatic preparation, causing a notable fidelity drop that can only be mitigated by slowing the parameter variation rate. The following section will therefore test this state preparation for a variety of variation rates across stages 2 and 3, while also exposing them to phase noise.
@@ III.2 Introducing phase noise
When noise is added to the transverse field, the resulting dephasing of the dynamics accumulates and, over time, causes unwanted excitations. Noise is added to the system with $\phi(t)$, which is implemented as a phase term modulating $\Omega(t)$ in Eq. (II). Realizations of $\phi(t)$ are generated using an adaptation of the TK95 algorithm [61] that stochastically samples a power spectrum chosen to be typical for Rydberg lasers. Figure 3(a) provides the original frequency power spectrum of laser noise for a diode laser, and the corresponding phase power spectrum that can then be used to generate the unique discrete noise signals shown in Figure 3(b). Details on the simulation of realistic phase noise and the excitation dynamics that lead to the results in this paper are provided in the Appendices A and C.
Figure 4 shows how different adiabatic evolution times result in a drop in fidelity of the final prepared state with the instantaneous ground state at the end of stage 2 in plot (a), and at the end of stage 3 in plot (c) which marks the end of the adiabatic protocol. Note that the evolution time of stage 1 is fixed to $T_{1}V_{dd}=5$ as this is sufficient to preserve high fidelity. To generate data for stage 3 we use the prepared state at the end of stage 2 that corresponds to the highest fidelity. In addition, the energy of the final prepared state relative to the ground sate is also provided, in (b) for stages 1 and 2, and in (d) for stage 3. Each data point in Figure 4 thus represents a final fidelity or energy at the end of adiabatic ramps simulated for a separate evolution time and system size, averaged across 100 independent noise realizations of phase noise, while a black line corresponds to results for an 11-site system without laser phase noise for reference. When no laser noise is present, we see continued improvement in fidelity as the evolution time increases and diabatic excitations are suppressed. This improvement is rapid for short evolution times and more gradual in the case of longer ramp times. However, the addition of phase noise causes an overall drop in fidelity for preparation across all evolution times, and in particular for longer evolution times where the cumulative dephasing leads to prolonged excitation out of the ground state. This difference in fidelity and energy is present for both stages and is best observed for the system size of $N=11$, with a clear divergence between realizations with and without noise.
Looking at the final fidelity of the prepared state at the end of stage 2, Figure 4(a) shows the competition between diabatic and dephasing excitations, since both have very different excitation profiles and opposite relationships to evolution time. While diabatic excitation involves low energy eigenstates and occurs for short evolution times after loss of adiabaticity during fast $\Omega$ and $\delta$ variation rates, dephasing excitation can affect higher energy states and occurs after the gradual accumulation of phase errors in the laser drive affecting longer evolution times. This gradual accumulation is precisely why the standard error in Figure 4 increases with time and individual trajectories begin to diverge as independent phase errors build up. The competition between the two excitation mechanisms leads to the emergence of an optimal time $T_{2O}$ (marked by magenta dots in Figure 4(a)), which is long enough to suppress most diabatic excitation but short enough to allow only small dephasing, thus corresponding to the largest achievable fidelity for a given system size. For stage 2 of the adiabatic state preparation $T_{2O}$ remains roughly consistent between different system sizes, as diabatic excitation remains low except for the shortest evolution times, and the frequency profile of the added noise is the same across all simulations, meaning that the generated noise signals take approximately the same time to begin affecting dynamics. Figure 4(b) shows the transition from mainly diabatic to mainly dephasing excitation in stage 2, which occurs at a comparatively short evolution time of $T_{2}V_{dd}\approx 60$ for all system sizes, as the ground state energy gap during this ramp remains large enough to permit faster variation rates without inducing excessive diabatic excitation. For stage 2 ramps that are longer than $T_{2}V_{dd}\approx 60$ diabatic excitation becomes insignificant, and we see a steady accumulation of energy from phase noise that is similar across all system sizes since the same noise profile is being used. Even though the same phase noise energy is added into the system, fidelities drop more for larger system sizes which have a larger density of states and smaller energy gaps leading to more excitation out of the ground state. In the case of energy, we see from Figures 4(b) that the differences in the amount of energy added to the system between different system sizes are mainly governed by diabatic excitation. Since diabatic excitation across the first two stages is minimal due to a large ground state energy gap, differences between energies only occur for very short evolution times where diabatic excitation is the larger for larger systems since they have a higher density of states and smaller energy gaps overall. For longer adiabatic ramps we see energy rise linearly as a function of the time that phase noise has to accumulate. Moving on to stage 3 in Figures 4(c,d), we see that the dynamics become very different due to the dramatic increase in diabatic excitation as the system passes through a critical region in the parameter space where the ground state energy gap becomes very small. In Figure 4(c) this increase in diabatic excitation leads to much larger losses in fidelity that are more acute for larger system sizes, as they exhibit a larger density of states and smaller energy gaps. The improvements in fidelity from mitigating diabatic excitation when moving from short to longer evolution times are also much more gradual in stage 3 than across stages 1 and 2 where ramps shorter than $T_{2}V_{dd}=50$ were strongly affected and the optimal ramp time (which signals a transition from diabatic to dephasing excitation) remained consistently around $T_{2}V_{dd}\approx 75$. Instead, in stage 3 we observe that the optimal evolution times increase with system size, from $T_{3}V_{dd}=100$ for 11 sites, to $T_{3}V_{dd}=167$ for 17 sites, and $T_{3}V_{dd}=200$ for 23 sites, with the 31 site system yet to reach its zenith after a longest simulated time of $T_{3}V_{dd}=250$. Larger systems sizes are increasingly more susceptible to diabatic excitations and the variation rate must reduce considerably even at the cost of more dephasing. This is consistent with recent analysis [46], predicting exceedingly long adiabatic ramps for larger system sizes due to diabatic losses.
Figure 5: Final fidelities of the prepared state $|\psi(T_{3})\rangle$ with $Z_{2}$ ordered ground state $|\psi_{gr}(T_{3})\rangle$ for the adiabatic state preparation described in Figure 1 with optimal evolution times for stage 2 ($T_{2}V_{dd}=82$) and stage 3 ($T_{3}V_{dd}=83$), and two varying parameters. (a) Final fidelity for different linear scalings of magnitude $m$ are applied directly to phase noise $\phi(t)\rightarrow m\phi(t)$. (b) Final fidelities for different values of $V_{dd}/2\pi$ ranging from 0.1100.1-10MHz. Colored lines show results for different system sizes. Error bars show standard error after averaging 100 preparations with unique laser phase noise signals. These results have been calculated using tensor network simulations using the iTensors package in Julia [18], with time evolution achieved using the TDVP algorithm with a maximum bond dimension of 100.
Continuing the analysis, Figure 5 demonstrates the impact that changes in $V_{dd}$ and noise strength have on the performance of the adiabatic protocol for different system sizes while running with optimal variation rates. To begin with, Figure 5(a) demonstrates how the protocol performs for different noise strengths by introducing a factor $m$ that linearly scales the phase noise to $\phi(t)\rightarrow m\phi(t)$. The susceptibility of such adiabatic protocols to phase noise is clear, as we observe an almost complete degradation in fidelity with the target ground state within a single order of magnitude of amplification. Figure 5(b) shows the final fidelities after performing the protocol for a variety of $V_{dd}$ values which, given that the laser parameters used are expressed in units of $V_{dd}$, shift the energy scaling of the entire system. Since the frequency profile of the noise remains constant with a broad peak in power at 480480kHz (see Figure 3), tuning $V_{dd}$ directly shifts the transitions to higher energy levels in and out of resonance. Phase noise made up of frequencies that are too high or too low relative to the dynamics of the system inhibits the transfer of power into the system, thereby causing less excitation out of the ground state. This is why Figure 5(b) shows that the lowest energy added to the system, and thus the highest fidelity achieved, occurs as far as possible from the peak noise frequency at $V_{dd}=10$MHz, with a dramatic degradation in fidelity as the system dynamics move towards affected frequencies at $V_{dd}=0.5$MHz, and a visible improvement as they are surpassed. The result motivates the choice of 1010MHz as a reasonable regime to test the performance of the protocol in Figure 4, and demonstrates that even if phase noise is present in a laser, the effects of phase noise on many-body experiments can be effectively mitigated by moving dynamics far away from affected frequencies. A similar conclusion was previously shown for single atoms and two qubit gates in [12, 39].
Studying the effects of phase noise on fidelity provides useful experiment-specific insight on imperfect adiabatic ramps in many-body systems. However, the measure considers only the ground state, revealing little about how dephasing distributes excitations across the energy eigenstates of the system. To get a deeper understanding of the mechanisms through which dephasing excitation propagates we have to analyze the full Hilbert space. The following section provides an analysis of the excitation dynamics for laser phase noise by studying the full energy spectrum of a one-dimensional chain of Rydberg atoms.